Haley Photocopying purchases paper from an​ out-of-state vendor. Average weekly demand for paper is 110 cartons per week for which Haley pays ​$20 per carton. Inbound shipments from the vendor average 950 cartons with an average lead time of 5 weeks. Haley operates 52 weeks per​ year; it carries a 6 week supply of inventory as safety stock and no anticipation inventory. The vendor has recently announced that they will be building a facility near Haley Photocopying that will reduce lead time to 2 week. ​Further, they will be able to reduce shipments to 150 cartons. Haley believes that they will be able to reduce safety stock to a 2 week supply. What impact will these changes make to​ Haley's average inventory level and its average aggregate inventory​ value?

The changes decrease​ Haley's average aggregate inventory level by

cartons. ​(Enter your response as a whole​ number.)

The changes decrease​ Haley's average aggregate inventory value by

​(Enter your response as a whole​ number.)
Answer #1

Initial case

Average inventory = Safety stock + Q/2

Avg inventory = 6*110+950/2 = 1135

Value = 1135*20= 22700

With near facility

Avg inventory= 2*110+150/2 = 295

Value = 295*20 = 5900

Change in inventory level 840

Change in value 16800